# Astrophysical Black holes: An Explanation for the Galaxy Quenching

### v2

@@ V Summary and Conclusions
The 2017 EHT observations of Sagittarius A*, identified an ABH without an event horizon (EH), i.e., JMN1, is one of the best possible black hole mimickers for Sgr A*[89]. Considering the various aspects of these models, we have proposed here that the ABH scenario offers a more plausible explanation for the sustainable quenching of the galaxy. The two toy models, namely JMN1 & JMN2, have been proposed as the globally naked singularities, which form through the gravitational collapse of the physically relevant matter clouds. These models are just examples among the larger group of equilibrium configurations that could arise from the gravitational collapse, throughout which the metric of the collapsing cloud assumes the forms given in Eqn. (1) & (2). The existence of the event horizon around the astrophysical black holes has been under question since the inception of the cosmic censorship conjecture(CCC). The gravitational collapse of the matter clouds could end in various possible entities, including black holes, ultra-dense compact objects without event horizons, globally and locally naked singularities, etc. The black holes in the AGNs have always been central to explaining the quenching of the galaxy and host of other astrophysical manifestations. From Silk and Rees in 1998 to the present times, the role of BHs in galaxy quenching has been explored through various mathematical tools and simulations[90, 28, 41]. The various shortcomings of the AGN feedback models are noted in Sect. 3, which provided us the impetus to explore the role of some alternative end-products of the gravitational collapse of the matter clouds, with apparently more realizable physical conditions.
@@
We have also briefly discussed the possibility of observing the StMABHs, which can produce energies many powers of ten above the Eddington rate, until they get enveloped by an event horizon. The Parker winds, having powerful electric currents, could signal this transition. This paper is a preliminary attempt to outline possible scenarios which can be tested observationally, as well as simulated for various accretion disk models, as already done for the BHs[91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110]. The key difference in simulating the ABHs without EHs and the normal BHs would be in terms of radiative efficiency which is linked to the ISCO. It can therefore be expected that, for a given mass, ABHs would be more effective in galaxy quenching, compared to the normal BHs,

### v3

@@ V Summary and Conclusions
The 2017 Event Horizon Telescope (EHT) observations of Sagittarius $A^{*}$ identified an ABH without a traditional event horizon, i.e., the JMN1 spacetime represents one of the best black hole mimickers for Sgr $A^{*}$ [61]. As stated, “More strikingly, for the same approximate luminosity as Sgr A* at 200 GHz, accretion flows onto these singularities have spectra nearly identical to those of a Schwarzschild black hole, indicating that a JMN-1 naked singularity with a photon sphere may be one of the best possible black hole mimickers for Sgr $A^{*}$[61]. Considering the various aspects of these models, we have proposed here that the ABH scenario offers a more plausible explanation for the sustainable quenching of the galaxy. The two toy models, namely JMN1 & JMN2, have been proposed as the globally naked singularities, which form through the gravitational collapse of the physically relevant matter clouds. These models are just examples among the larger group of equilibrium configurations that could arise from the gravitational collapse, throughout which the metric of the collapsing cloud assumes the forms given in Eqn. (1) & (2). The existence of the event horizon around the astrophysical black holes has been under question since the inception of the cosmic censorship conjecture(CCC). The gravitational collapse of the matter clouds could end in various possible entities, including black holes, ultra-dense compact objects without event horizons, globally and locally naked singularities, etc. The black holes in the AGNs have always been central to explaining the quenching of the galaxy and host of other astrophysical manifestations. From Silk and Rees in 1998 to the present times, the role of BHs in galaxy quenching has been explored through various mathematical tools and simulations[104, 28, 41]. The various shortcomings of the AGN feedback models are noted in Sect. 3, which provided us the impetus to explore the role of some alternative end-products of the gravitational collapse of the matter clouds, with apparently more realizable physical conditions.
@@
We have also briefly discussed the possibility of observing the StMABHs, which can produce energies many powers of ten above the Eddington rate, until they get enveloped by an event horizon. The Parker winds, having powerful electric currents, could signal this transition. This paper is a preliminary attempt to outline possible scenarios which can be tested observationally, as well as simulated for various accretion disk models, as already done for the BHs[105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124]. The key difference in simulating the ABHs without EHs and the normal BHs would be in terms of radiative efficiency which is linked to the ISCO. It can therefore be expected that, for a given mass, ABHs would be more effective in galaxy quenching, compared to the normal BHs.
We would like to highlight several future directions motivated by this work. Active Galactic Nuclei (AGNs) are conventionally modeled with supermassive black holes. In this study, we propose an alternative possibility involving horizonless ultra-compact objects, referred to here as Astrophysical Black Holes (ABHs), which may arise from the gravitational collapse of physically relevant matter clouds[60, 81]. These objects could offer a different perspective on long-term quenching of star formation. A natural next step is the explicit modeling of AGNs powered by ABHs without an event horizon. Current observations, including the Event Horizon Telescope image of Sgr A, do not yet allow a definitive distinction between black holes and certain naked singularity configurations. Developing AGN models based on ABHs and comparing their predicted galactic effects with those of standard black hole models may therefore provide new insights into fundamental questions such as the physical existence of event horizons and the validity of the cosmic censorship conjecture. If supported by future studies, ABH-based scenarios may also help explain long-term quenching and could imply an alternative pathway for galaxy evolution. In the context of stellar-mass ABHs[44, 45], we have suggested potential observational signatures that may appear prior to horizon formation. Further investigation along these lines could improve our understanding of the late stages of continual gravitational collapse, a regime characterized by ultra-strong gravity where quantum effects are expected to become important. This aspect becomes even more important in the context of recent JWST observations reporting a compact quenched galaxy at $z=7.3$, when the universe was only about 700 million years old. Explaining such early quenching generally requires a preceding starburst event, followed by strong feedback that suppresses star formation[77]. In this scenario, stellar-mass Astrophysical Black Holes (StMABHs) could provide the necessary feedback. Further exploration of this possibility may offer insight into quenching mechanisms in the early universe and could also shed light on the potential formation of naked singularities during epochs when the earliest galaxies became visible. Finally, if magnetic field structures in outflows — particularly winds resembling Parker-type solutions and carrying strong electric currents — are interpreted as signatures of matter or radiation emerging from naked singularities, their detailed study may offer a phenomenological route toward probing quantum gravity effects. Such observationally guided approaches could help bridge the gap between astrophysical measurements and current theoretical models of quantum gravity.